OpenStudy (anonymous):

PLEAASSSSE HELP!!!! 1!!!!!!!!!!!!!!!!!!!!!! Suppose a triangle has sides a b c, and let theta be the angle opposite the side length a. if cos theta>0, what must be true a) a^2+b^2=c^2 b)a^2+b^2>c^2 c) b^2+c^2>a^2 d)b^2+c^2

OpenStudy (here_to_help15):

We have a law of cosines which says cos A = (b^2 + c^2 - a^2)/2bc cos B = (a^2 + c^2 - b^2)/2ac cos C = (a^2 + b^2 - c^2)/2ab where A,B,C are angles opp. to sides a,b and c so your ques. cos theta >0 (a^2 + c^2 - b^2)/2ac > 0 solving you'll get a^2 + c^2 > b^2 its b

OpenStudy (anonymous):

thank u so much i hope this is right!!!:)

OpenStudy (here_to_help15):

k : )